Solve
(D-3D'-2)^2.z= 2e^2x.sin(2y+x)
1
Expert's answer
2020-11-19T15:16:04-0500

If equation is of the form, "(D-mD'-k)z =0" has the solution "C.F. = e^{kx}\phi_1(y+mx)"


Then, for the given equation,

"(D-3D'-2)^2z= 2e^{2x}sin(2y+x)"


C.F. of the equation, "C.F. = e^{2x}\phi_1(y+3x)+xe^{2x}\phi_2(y+3x)"


P.I. is, "\frac{1}{[D-3D'-2]^2}2e^{2x}sin(2y+x) = 2e^{2x}\frac{1}{[(D+2)-3D'-2]^2}sin(2y+x)"


"=2e^{2x}\frac{x^2}{2!}sin(2y+x)={x^2}e^{2x}sin(2y+x)"


Then solution is,

"z = e^{2x}\phi_1(y+3x)+xe^{2x}\phi_2(y+3x)+{x^2}e^{2x}sin(2y+x)"